A sample of 100 workers in a large plant gave a mean assembly time of 294 seconds with a standard deviation of 12 seconds in a time and motion study. Find a 95%percent confidence interval for the mean assembly time for all the workers in the plant

1
Expert's answer
2021-12-14T05:43:33-0500

The confidence interval for population mean can be estimated the next way:

"x-Cr_a*{\frac s {\sqrt{n}}}<u<x+Cr*{\frac s {\sqrt{n}}}" , where x - sample mean, "Cr_a" - critical value at a confidence level, s - sample standard deviation, n - sample size, u - population mean

Since the sample size is big, we can use z-score(standard normal distribution) as critical value, so

"P(N(0,1)<Cr)=1-{\frac a 2}=0.975\implies Cr=1.96"

So, we have

"294-1.96*{\frac {12} {\sqrt{1000}}}<u<294+1.96*{\frac {12} {\sqrt{1000}}}\implies293.26<u<294.74"